# Generalization Error of the Tilted Empirical Risk

### v2

@@ Abstract
The generalization error (risk) of a supervised statistical learning algorithm quantifies its prediction ability on previously unseen data. Inspired by exponential tilting, Li et al. (2021) proposed the tilted empirical risk as a non-linear risk metric for machine learning applications such as classification and regression problems. In this work, we examine the generalization error of the tilted empirical risk. In particular, we provide uniform and information-theoretic bounds on the tilted generalization error, defined as the difference between the population risk and the tilted empirical risk, with a convergence rate of $O(1/\sqrt{n})$ where $n$ is the number of training samples. Furthermore, we study the solution to the KL-regularized expected tilted empirical risk minimization problem and derive an upper bound on the expected tilted generalization error with a convergence rate of $O(1/n)$.
@@ 1 Introduction
Empirical risk minimization (ERM) is a popular framework in machine learning. The performance of the empirical risk (ER) is affected when the data set is strongly imbalanced or contains outliers. For these scenarios, inspired by the log-sum-exponential operator with applications in multinomial linear regression and naive Bayes classifiers (Calafiore et al., 2019; Murphy, 2012; Williams and Barber, 1998), the tilted empirical risk (TER) is proposed by Li et al. (2021) for supervised learning application, such as classification and regression problems. Li et al. (2021, 2023a) showed that tilted empirical risk minimization (TERM) can handle class imbalance, mitigate the effect of outliers, and enable fairness between subgroups. Different applications of TERM have been explored, e.g., differential privacy (Lowy and Razaviyayn, 2021), semantic segmentation (Szabó et al., 2021) and noisy label self-correction (Zhou et al., 2020). In this paper, we aim to corroborate the empirical success of the TERM framework through statistical learning theory.
@@
A central concern in statistical learning theory is understanding the efficacy of a learning algorithm when applied to test data. This evaluation is typically carried out by investigating the generalization error, which quantifies the disparity between the performance of the algorithm on the training dataset and its performance on previously unseen data, drawn from the same underlying distribution, via a risk function. Understanding the generalization behaviour of learning algorithms is one of the most important objectives in statistical learning theory. Various approaches have been developed (Rodrigues and Eldar, 2021), including VC dimension-based bounds (Vapnik, 1999), stability-based bounds (Bousquet and Elisseeff, 2002b), PAC Bayesian bounds (McAllester, 2003), and information-theoretic bounds (Russo and Zou, 2019; Xu and Raginsky, 2017). However, the TER generalization performance under unbounded loss function has not yet been studied. This paper focuses on the generalization error of the tilted empirical risk (tilted generalization error) of learning algorithms. Our contributions here can be summarized as follows,
  • We provide upper and lower bounds on the tilted generalization error via uniform and information-theoretic approaches under bounded loss functions for both positive and negative tilts and show that the convergence rates of the upper bounds are $O(1/\sqrt{n})$, as expected, where $n$ is the number of training samples.
@@
  • We provide upper and lower bounds on the tilted generalization error under unbounded loss functions for the negative tilt with the convergence rate of $O(1/\sqrt{n})$ via a uniform approach.
@@
  • We study the robustness of the tilted empirical risk under distribution shift induced by noise or outliers for unbounded loss functions with bounded second-moment assumption and negative tilt, and derive generalization bounds that justify the robustness properties of TERM (Li et al., 2021) for negative tilt.
@@
  • We study the KL-regularized TERM problem and provide an upper bound on the expected tilted generalization error with convergence rate $O(1/n)$.
The paper is organised as follows: Section 2 introduces notation, the problem, and the risk functions used in this paper. Our upper and lower bounds on the tilted generalization error for bounded loss functions via uniform and information-theoretic approaches are given in Section 3. The upper and lower bounds on the tilted generalization error for unbounded loss functions with bounded second moments are given in Section 4. Section 5 is devoted to the study of robustness to the distributional shift in training samples. The KL-regularized TERM is considered in Section 6. Section 7 surveys related work. Section 8 concludes the paper. Technical tools and proofs are deferred to the appendices.
@@ 2.1 Problem Formulation
Let $S=\{Z_{i}\}_{i=1}^{n}$ be the training set, where each sample $Z_{i}=(X_{i},Y_{i})$ belongs to the instance space $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$; here $\mathcal{X}$ is the input (feature) space and $\mathcal{Y}$ is the output (label) space. We assume that $Z_{i}$ are i.i.d. generated from the same data-generating distribution $\mu$.
Here we consider the set of hypothesis $\mathcal{H}$ with elements $h:\mathcal{X}\mapsto\mathcal{Y}\in\mathcal{H}$. When $\mathcal{H}$ is finite, then its cardinality is denoted by $\mathrm{card}(\mathcal{H})$. In order to measure the performance of the hypothesis $h$, we consider a non-negative loss function :×$\ell:\mathcal{H}\times\mathcal{Z}\to\mathbb{R}_{0}^{+}$.
We apply different methods to study the performance of our algorithms, including uniform and information-theoretic approaches. In uniform approaches, such as the VC-dimension and the Rademacher complexity approach (Vapnik, 1999; Bartlett and Mendelson, 2002), the hypothesis space is independent of the learning algorithm. Therefore, these methods are algorithm-independent; our results for these methods do not specify the learning algorithms.
Learning Algorithms: For information-theoretic approaches in supervised learning, following Xu and Raginsky (2017), we consider learning algorithms that are characterized by a Markov kernel (a conditional distribution) $P_{H|S}$. Such a learning algorithm maps a data set $S$ to a hypothesis in $\mathcal{H}$, which is chosen according to $P_{H|S}$. This concept thus includes randomized learning algorithms.
Robustness: Suppose that due to an outlier or noise in training samples, the underlying distribution of the training dataset, $\hat{S}$ is shifted via the distribution of noise or outlier in the training dataset, denoted as $\tilde{\mu}\in\mathcal{P}(\mathcal{Z})$. We model the distributional shift via distribution $\tilde{\mu}$ due to inspiration by the notion of influence function (Marceau and Rioux, 2001; Christmann and Steinwart, 2004; Ronchetti and Huber, 2009).
@@ 2.2 Risk Functions
$\mathrm{gen}(h,S):=\mathrm{R}(h,\mu)-\widehat{\mathrm{R}}(h,S);$ (2)
this is the difference between the true risk and the linear empirical risk. The TER, as a non-linear empirical risk (Li et al., 2021) and estimator of population risk, is defined by
@@
The TER is an increasing function in $\gamma$ (Li et al., 2023a, Theorem 1), and as $\gamma\rightarrow 0$, the TER converges to the linear empirical risk in (1). Inspired by (Li et al., 2021), the primary objective is to optimize the population risk; the TERM is utilized in order to help the learning dynamics. Therefore, we decompose the population risk as follows:
@@
where $g_{t}$ is a real function. Furthermore, we denote the excess risk under the tilted emprical risk as,
$\mathfrak{E}_{\gamma}(\mu):=\mathrm{R}(h_{\gamma}^{*}(S),\mu)-\mathrm{R}(h^{*}(\mu),\mu),$ (7)
where $h^{*}(\mu):=\arg\min_{h\in\mathcal{H}}\mathrm{R}(h,\mu)$ and $h_{\gamma}^{*}(S):=\arg\min_{h\in\mathcal{H}}\hat{\mathrm{R}}_{\gamma}(h,S)$.
We denote the expected TER with respect to the distribution of $S$ by
$\begin{split}\overline{\mathrm{R}}_{\gamma}(h,P_{S})=\mathbb{E}_{P_{S}}[\widehat{\mathrm{R}}_{\gamma}(h,S)],\end{split}$ (8)
and we denote the tilted (true) population risk by
$\begin{split}\mathrm{R}_{\gamma}(h,P_{S})=\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{P_{S}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))\Big{]}\Big{)}.\end{split}$ (9)
Under the i.i.d. assumption, the tilted population risk is equal to an entropic risk function (Howard and Matheson, 1972). We also introduce the non-linear generalization error, which plays an important role in obtaining our bounds, as
$\widehat{\mathrm{gen}}_{\gamma}(h,S):=\mathrm{R}_{\gamma}(h,\mu)-\widehat{\mathrm{R}}_{\gamma}(h,S).$ (10)
Information-theoretic Approach: For the information-theoretic approach, as the hypothesis $H$ is a random variable under a learning algorithm as Markov kernel, i.e., $P_{H|S}$, we take expectations over the hypothesis $H$ over the above expressions for fixed $h$ to define the expected true risk, expected empirical risk, and expected tilted generalization error,
$\begin{split}\mathrm{R}(H,P_{H}\otimes\mu)&:=\mathbb{E}_{P_{H}\otimes\mu}[\ell(H,Z)],\\ \overline{\mathrm{R}}_{\gamma}(H,P_{H,S})&:=\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)],\\ \overline{\mathrm{gen}}_{\gamma}(H,S)&:=\mathbb{E}_{P_{H,S}}[\mathrm{gen}_{\gamma}(H,S)].\end{split}$ (11)
Similar to (11), we define the tilted population risk and the tilted generalization error, (9) and (10), as expectations. We also provide upper bounds on the absolute value of expected tilted generalization error with respect to the joint distribution of $S$ and $H$, of the form
@@
where $g_{e}$ is a real function.
3 Generalization Bounds for Bounded Loss Function
Upper bounds under linear empirical risk for bounded loss functions via information theoretic and uniform bounds are studied by Shalev-Shwartz and Ben-David (2014) and Xu and Mannor (2012), respectively. Inspired by these works, in this section, we provide upper bounds on the tilted generalization error via uniform and information-theoretic approaches for bounded loss functions with the convergence rate of $O(1/\sqrt{n})$ which is similar to generalization error under linear empirical risk. Upper bounds via stability (Bousquet and Elisseeff, 2002b), Rademacher complexity (Bartlett and Mendelson, 2002) and PAC-Bayesian approaches (Alquier, 2021) are provided in Appendix H. All proof details are deferred to Appendix D.
In this section the following assumption is made.
Assumption 3.1 (Bounded loss function).
Assumption 3.1 will be relaxed in Section 4.
@@ 3.1 Uniform Bounds
For uniform bounds of the type (6) we decompose the tilted generalization error (4) as follows,
$\begin{split}&\mathrm{gen}_{\gamma}(h,S)=\underbrace{\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})}_{I_{1}}\\ &\qquad+\underbrace{\mathrm{R}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S)}_{I_{2}},\end{split}$ (12)
where $I_{1}$ is the difference between the population risk and the tilted population risk and $I_{2}$ is the non-linear generalization error.
@@
Under Assumption 3.1, for $\gamma\in\mathbb{R}$ the difference between the population risk and the tilted population risk satisfies
$\begin{split}&\frac{-1}{2\gamma}\mathrm{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}\leq\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})\\ &\leq\frac{-\exp(-2\gamma M)}{2\gamma}\mathrm{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}.\end{split}$ (13)
Note that for $\gamma\rightarrow 0$, the upper and lower bounds in Proposition 3.2 are zero.
As the log function is Lipschitz on a bounded interval, applying the Hoeffding inequality to term $I_{2}$ and Proposition 3.2 to term $I_{1}$ in (12), we obtain the following upper bound on the tilted generalization error.
@@
Given Assumption 3.1, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$ the tilted generalization error satisfies the upper bound,
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\leq\frac{-\exp(-2\gamma M)}{2\gamma}\operatorname{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}\\ &\qquad+\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.\end{split}$
Theorem 3.4.
Under the same assumptions of Theorem 3.3, for a fixed $h\in\mathcal{H}$, with probability at least $(1-\delta)$, the tilted generalization error satisfies the lower bound
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\geq\frac{-1}{2\gamma}\operatorname{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}\\ &\qquad-\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.\end{split}$ (14)
Remark 3.5.
The lower bound in Theorem 3.4 for a negative $\gamma$ can be tighter than the lower bound on the tilted generalization error for a positive $\gamma$, (14), as the variance is positive.
Combining Theorem 3.3 and Theorem 3.4, we derive an upper bound on the absolute value of the titled generalization error.
Corollary 3.6.

### v3

# Generalization and Robustness of the Tilted Empirical Risk

@@ Abstract
The generalization error (risk) of a supervised statistical learning algorithm quantifies its prediction ability on previously unseen data. Inspired by exponential tilting, Li et al. (2021) proposed the tilted empirical risk (TER) as a non-linear risk metric for machine learning applications such as classification and regression problems. In this work, we examine the generalization error of the tilted empirical risk in the robustness regime under negative tilt. Our first contribution is to provide uniform and information-theoretic bounds on the tilted generalization error, defined as the difference between the population risk and the tilted empirical risk, under negative tilt for unbounded loss function under bounded $(1+\epsilon)$-th moment of loss function for some $\epsilon\in(0,1]$ with a convergence rate of $O(n^{-\epsilon/(1+\epsilon)})$ where $n$ is the number of training samples, revealing a novel application for TER under no distribution shift. Secondly, we study the robustness of the tilted empirical risk with respect to noisy outliers at training time and provide theoretical guarantees under distribution shift for the tilted empirical risk. We empirically corroborate our findings in simple experimental setups where we evaluate our bounds to select the value of tilt in a data-driven manner.
Tilted Empirical Risk, Generalization Error, Robustness, Excess Risk
[…]
@@ 1 Introduction
Empirical risk minimization (ERM) is a popular framework in machine learning. The performance of the empirical risk (ER) is affected when the data set is strongly imbalanced or contains outliers. For these scenarios, inspired by the log-sum-exponential operator with applications in multinomial linear regression and naive Bayes classifiers (Calafiore et al., 2019; Murphy, 2012; Williams & Barber, 1998), Li et al. (2021) propose the tilted empirical risk (TER) for supervised learning applications, such as classification and regression problems. Li et al. (2021, 2023a) showed that tilted empirical risk minimization (TERM) can handle class imbalance, mitigate the effect of outliers, and enable fairness between subgroups. Different applications of TERM have been explored, e.g., differential privacy (Lowy & Razaviyayn, 2021), semantic segmentation (Szabó et al., 2021), noisy label self-correction (Zhou et al., 2020), and off-policy learning and evaluation (Behnamnia et al., 2025). [^1] Behnamnia et al. (2025) use the term log-sum-exponential, which is the same as the tilted empirical risk studied herein. In this paper, we corroborate the empirical success of the TERM framework through statistical learning theory.
@@
A central concern in statistical learning theory is understanding the efficacy of a learning algorithm when applied to test data. This evaluation is typically carried out by investigating the generalization error, which quantifies the disparity between the performance of the algorithm on the training dataset and its performance on previously unseen data, drawn from the same underlying distribution, via a risk function. Understanding the generalization behaviour of learning algorithms is one of the most important objectives in statistical learning theory. Various approaches have been developed for empirical risk (Rodrigues & Eldar, 2021), including VC dimension-based bounds (Vapnik, 1999), stability-based bounds (Bousquet & Elisseeff, 2002b), PAC Bayesian bounds (McAllester, 2003), and information-theoretic bounds (Russo & Zou, 2019; Xu & Raginsky, 2017). This paper focuses on the generalization and robustness of the tilted empirical risk (tilted generalization error) of learning algorithms. Our contributions are:
@@
  • In Section 3, we provide upper and lower bounds on the tilted generalization error under unbounded loss functions for the negative tilt with a bound via uniform and information-theoretical approaches and establish the convergence rate of $O(n^{-\epsilon/(1+\epsilon)})$ for some $\epsilon\in(0,1]$.
@@
  • In Section 4, we study the robustness of the tilted empirical risk under distribution shift induced by noise or outliers for unbounded loss functions with bounded $(1+\epsilon)$-th moment assumption, for some $\epsilon\in(0,1]$, and negative tilt, and derive generalization bounds.
  • In Section 5, we provide a data-driven approach to selecting the value of tilt for regression problems, with observed test errors corroborating the theoretical bounds.
@@
  • In Section 6, we study the KL-regularized TERM problem under unbounded loss function with bounded second moment assumption and provide an upper bound on the expected tilted generalization error with convergence rate $O(n^{-\epsilon})$ under unbounded loss function with bounded $(1+\epsilon)$-th moment for some $\epsilon\in(0,1]$.
@@ 2.1 Problem Formulation
Let $S=\{Z_{i}\}_{i=1}^{n}$ be the training set, where each sample $Z_{i}=(X_{i},Y_{i})$ belongs to the instance space $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$; here $\mathcal{X}$ is the input (feature) space and $\mathcal{Y}$ is the output (label) space. We assume that $Z_{i}$ are i.i.d. generated from the same data-generating distribution $\mu$. We also assume $\tilde{Z}\sim\mu$ as an i.i.d. sample with respect to the training set.
A set of hypotheses $\mathcal{H}$ has elements $h:\mathcal{X}\mapsto\mathcal{Y}\in\mathcal{H}$. When $\mathcal{H}$ is finite, then its cardinality is denoted by $\mathrm{card}(\mathcal{H})$. In order to measure the performance of the hypothesis $h$, we use a non-negative loss function :×$\ell:\mathcal{H}\times\mathcal{Z}\to\mathbb{R}_{0}^{+}$.
We apply different methods to study the performance of our algorithms, including uniform and information-theoretic approaches. In uniform approaches, such as the VC-dimension and the Rademacher complexity approach (Vapnik, 1999; Bartlett & Mendelson, 2002), the hypothesis space $\mathcal{H}$ is independent of the learning algorithm. Therefore, these methods are algorithm-independent; our results for these methods do not specify the learning algorithms.
Learning Algorithms: For information-theoretic approaches in supervised learning, following Xu & Raginsky (2017), we consider learning algorithms that are characterized by a Markov kernel (a conditional distribution) $P_{H|S}$. Such a learning algorithm maps a data set $S$ to a hypothesis in $\mathcal{H}$, which is chosen according to $P_{H|S}$. This concept thus includes randomized learning algorithms.
@@ 2.2 Risk Functions
$\mathrm{gen}(h,S):=\mathrm{R}(h,\mu)-\widehat{\mathrm{R}}(h,S).$ (2)
This is the difference between the true risk and the linear empirical risk. The TER, as a non-linear empirical risk with tilt $\gamma$ (Li et al., 2021), a.k.a. log-sum-exponential, estimates the population risk by
@@
The TER is an increasing function in the tilt parameter $\gamma$ (Li et al., 2023a, Theorem 1), and as $\gamma\rightarrow 0$, the TER converges to the linear empirical risk in (1). In this work, we focus on negative tilt $\gamma<0$. Inspired by Li et al. (2021), the primary objective is to optimize the population risk; the TERM is used in order to help the learning dynamics. For our analysis, we decompose the population risk as follows:
@@
where $g_{t}$ is a real function. We set $h^{*}(\mu):=\arg\min_{h\in\mathcal{H}}\mathrm{R}(h,\mu)$ and $h_{\gamma}^{*}(S):=\arg\min_{h\in\mathcal{H}}\hat{\mathrm{R}}_{\gamma}(h,S)$. We use the following notations. The excess risk under the tilted empirical risk is
$\mathfrak{E}_{\gamma}(\mu):=\mathrm{R}(h_{\gamma}^{*}(S),\mu)-\mathrm{R}(h^{*}(\mu),\mu).$ (7)
The expected TER with respect to the distribution of $S$ is
$\begin{split}\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})=\mathbb{E}_{\mu^{\otimes n}}[\widehat{\mathrm{R}}_{\gamma}(h,S)],\end{split}$ (8)
and the tilted (true) population risk is
$\begin{split}&\mathrm{R}_{\gamma}(h,\mu^{\otimes n})\\ &\quad=\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{\mu^{\otimes n}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(h,Z_{i}))\Big{]}\Big{)}.\end{split}$ (9)
Under the i.i.d. assumption, the tilted population risk is equal to an entropic risk function (Howard & Matheson, 1972). We also introduce the non-linear generalization error [^2] We refer to this as non-linear generalization error since a non-linear transformation of the population risk, instead of the population risk, is used. , as
$\widehat{\mathrm{gen}}_{\gamma}(h,S):=\mathrm{R}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S).$ (10)
For ease of notation, we consider $\mathrm{R}_{\gamma}(h,\mu^{\otimes n})=\mathrm{R}_{\gamma}(h,\mu)$.
Information-theoretic Approach: For the information-theoretic approach, as the hypothesis $H$ is a random variable under a learning algorithm as Markov kernel, i.e., $P_{H|S}$, we take expectations over the hypothesis $H$. We denote the expected true risk, expected empirical risk, expected entropic risk, and expected tilted generalization error by
$\begin{split}&\mathrm{R}(H,P_{H}\otimes\mu):=\mathbb{E}_{P_{H}\otimes\mu}[\ell(H,Z)],\\ &\overline{\mathrm{R}}_{\gamma}(H,Q_{H,S}):=\mathbb{E}_{Q_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)],\\ &\mathrm{R}_{\gamma}(H,Q_{H,S}):=\\ &\qquad\frac{1}{\gamma}\log\Big{(}\mathbb{E}_{Q_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}\\ &\overline{\mathrm{gen}}_{\gamma}(H,S):=\mathbb{E}_{P_{H,S}}[\mathrm{gen}_{\gamma}(H,S)],\end{split}$ (11)
where $Q_{H,S}\in\{P_{H}\otimes\mu^{\otimes n},P_{H,S}\}$. In addition to bounds of the form (6), we provide upper bounds on the absolute value of expected tilted generalization error with respect to the joint distribution of $S$ and $H$, of the form
@@
where $g_{e}$ is a real function. We also introduce the non-linear expected generalization error, which plays an important role in deriving our bounds, as
$\widehat{\mathrm{gen}}_{\gamma}(H,S):=\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\mathrm{R}_{\gamma}(H,P_{H,S}).$ (12)
3 Generalization Bounds for Unbounded Loss Functions
In this section, we derive upper bounds on the tilted generalization error via uniform and information-theoretical approaches for negative tilt ($\gamma<0$) under bounded $(1+\epsilon)$-th moment of loss function, for some $\epsilon\in(0,1]$, with convergence rate of $O(n^{-\epsilon/(1+\epsilon)})$. Regarding the bounded loss function, we also provide the tilted generalization error bounds via uniform and information-theoretical approaches in Appendix H. Note that, with a bounded loss function, we can exploit the Lipschitz property of the logarithmic function. However, when dealing with unbounded losses, the loss function lacks the Lipschitz property, making it impossible to apply the same techniques used in the bounded case (see Appendix H).
Several works have already proposed some solutions to overcome the boundedness assumption under linear empirical risk (Haddouche & Guedj, 2022; Alquier & Guedj, 2018; Holland, 2019) via a PAC-Bayesian approach under the bounded second-moment assumption. Furthermore, an upper bound on generalization error via VC-dimension and growth function is proposed in Cortes et al. (2019, Corollary 12) with convergence rate of $O(\log(n)n^{-\epsilon/(1+\epsilon)})$. In contrast, we derive bounds with a convergence rate of $O(n^{-\epsilon/(1+\epsilon)})$. A more detailed comparison is provided in Section 7.
@@ 3.1 Uniform Bounds
The following assumption is made for the uniform analysis.
Assumption 3.1 (Uniform bounded ($1+\epsilon$)-th moment [^3] Note that we assume that higher order moments, larger than $(1+\epsilon)$, are unbounded.).
There is a constant $\kappa_{u}\in\mathbb{R}^{+}$ such that the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{\mu}[\ell^{1+\epsilon}(h,Z)]\leq\kappa_{u}^{1+\epsilon}$ uniformly for all $h\in\mathcal{H}$ and some $\epsilon\in(0,1]$.
The assumption on the $(1+\epsilon)$-th moment, Assumption 3.1, is satisfied for example if the loss function is sub-Gaussian or sub-Exponential under the distribution $\mu$ for all $h\in\mathcal{H}$, see Boucheron et al. (2013). Inspired by the approach of Behnamnia et al. (2025), which provides a regret bound based on a tilted operator, we aim to analyze the generalization error of tilted empirical risk within a uniform approach. All proof details for the results in this section are deferred to Appendix D.1.
For uniform bounds of the type (6), we decompose the tilted generalization error (4) as follows,
$\begin{split}&\mathrm{gen}_{\gamma}(h,S)=\underbrace{\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})}_{I_{1}}+\widehat{\mathrm{gen}}_{\gamma}(h,S),\end{split}$ (13)
@@
Under Assumption 3.1, for $\gamma<0$ and some $\epsilon\in(0,1]$, the difference between the population risk and the tilted population risk satisfies
$\begin{split}&0\leq\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})\leq|\gamma|^{\epsilon}\kappa_{u}^{1+\epsilon}.\end{split}$ (14)
Then, using Bernstein’s inequality (Boucheron et al., 2013) and properties of the logarithm, we can provide upper and lower bounds on the tilted generalization error.
Proposition 3.3.
Given Assumption 3.1 for some $\epsilon\in(0,1]$, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, then the following upper bound holds on the tilted generalization error for $\gamma<0$ and some $\epsilon\in(0,1]$,
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\leq\frac{2\exp(|\gamma|\kappa_{u})}{|\gamma|}\sqrt{\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{u}^{1+\epsilon}\log(2/\delta)}{n}}\\ &+\frac{4\exp(|\gamma|\kappa_{u})\log(2/\delta)}{3n|\gamma|}+|\gamma|^{\epsilon}\kappa_{u}^{1+\epsilon}.\end{split}$
Proposition 3.4.
Given Assumption 3.1 for some $\epsilon\in(0,1]$, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, and $\gamma<0$, the following lower bound on the tilted generalization error holds,
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\geq-\frac{2\exp(|\gamma|\kappa_{u})}{(1-\zeta)|\gamma|}\sqrt{\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{u}^{1+\epsilon}\log(2/\delta)}{n}}\\ &-\frac{4\exp(|\gamma|\kappa_{u})(\log(2/\delta))}{3n|\gamma|(1-\zeta)}.\end{split}$
Combining Proposition 3.3 and Proposition 3.4, we derive an upper bound on the absolute value of the tilted generalization error.